# Smooth Sailing: Lipschitz-Driven Uncertainty Quantification for Spatial Association

### v2

@@ 5 Discussion
We propose a new method for constructing confidence intervals for spatial linear models. We show via theory and experiments that our intervals accurately quantify uncertainty under misspecification and covariate shift. In experiments, our method is the only method that consistently achieves (or even comes close to) nominal coverage. We observe that, very commonly in spatial data analyses, covariates and responses may be observed at different locations in space. Since our method does not actually use the source covariate values in inference for $\theta^{\star}_{\text{OLS}}$, it can be applied in this common missing-data scenario. Though it requires additional work, we believe the ideas here will extend naturally to the widely used class of generalized linear models.
@@ Acknowledgments
This work was supported in part by an Office of Naval Research Early Career Grant, Generali, a Microsoft Trustworthy AI Grant, and NSF grant 2214177.
@@ Appendix A Extended Related Work
Bias in Spatial Regression
Linear models with a Gaussian process random effect — that is, models of the form
@@
with $g$ a Gaussian process and $\epsilon$ independent and identically distributed noise — are a classic tool in spatial regression and remain widely used in applications [69, 21, 24]. Whether to treat the covariates as fixed functions of spatial location or random is a topic of significant debate. In order for the model to be identifiable, it is generally necessary for the covariates to be thought of as random [19, Proposition 1]. However, Paciorek, [42] observed it may be more reasonable to think of covariates as fixed (the perspective we also take). If $X$ and $g$ are not independent (in the random case) or close to orthogonal in the fixed case, Paciorek, [42] illustrates that bias is introduced into the estimation of $\theta$. And that the degree of this bias depends on the relative scales over which $X$ and $g$ vary. He gives a closed form for this bias under strong linear-Gaussian assumptions that depend on parameters that would generally be unknown. Page et al., [43] builds on the result of Paciorek, [42], and shows that bias in predictions made by the spatial model due to confounding may be small, despite biases in parameter estimation. Hodges and Reich, [26] proposed using restricted spatial regression, essentially explaining as much of the variation in the response as possible by the covariates to reduce bias in estimation due to spatially-correlated error. We take this approach in the GLS baseline used in our experiments. Nobre et al., [41] extended earlier work about bias due to spatial confounding by considering the case when multiple observations are available at the same spatial location with independent noise. They showed that spatial regression models can still be biased with repeated observations due to confounding.
In our work, we focus on bias in estimation of the target conditional OLS. Because we do not assume the model is well-specified, there is no ‘true’ value of the parameter to estimate. Instead, our goal is to estimate the ‘best’ (in a least squares sense) linear approximation to the response. We focus on the case when $X$ is a fixed function of location (Assumption 1) and make weak non-parametric, smoothness assumptions (Assumption 4). This contrasts with prior works that generally assume the covariates are random and make linear-Gaussian assumptions to calculate the bias [42, 43]. Our work shows how to directly incorporate bias into confidence intervals for $\theta$ under our weak assumptions, but does not address issues of (non-)identifiability in the estimation problem.
@@
Local Versus Global Regression Estimates
An alternative perspective on regression in spatial settings, referred to as geographically weighted regression (GWR), focuses on estimating the association between the response and covariates at each location in space [6, 17]. For this regression problem to be well-defined, the covariates at each spatial location must be treated as random variables. Like our method, this approach uses nearest neighbors or local smoothing approaches to estimate the local coefficients at each location in space. And — as in our work — because of this “data-borrowing” from nearby locations, bias is introduced into the estimation of these local regression coefficients. The confidence intervals reported for GWR typically do not account for this bias [73]. While Yu et al., [73] provide a formula for the bias introduced by “data-borrowing”, their formula relies on the assumption that the model is well-specified and depends on unobservable parameters.
We focus on the estimation bias in the estimation of global parameters. Although local parameter estimates are often of interest in spatial analyses [17, Chapter 1], global estimates are also useful as summary statistics for making decisions that impact regions or multiple localities. And we characterize the bias in these estimates under weak, non-parametric assumptions and incorporate it into our confidence intervals. Extensions of our approach to GWR coefficients are a promising direction for future research. However, this extension is not simple, as it involves accounting for randomness in the covariates.
@@
In contrast, our approach ensures nominal coverage even when de-biasing the estimate via importance weighting is not possible. This advantage comes at the cost of an additional regularity assumption (Assumption 4), as well as often empirically wider confidence intervals.
@@ Appendix B Implementation Details
In this section, we describe the implementation details of our method. Particularly, this involves computing the upper bound on the bias, computation of $\Delta$ in Lemma 6 and computation of $\hat{\sigma}^{2}_{N}$ from Eq. 10. To summarize our method, we state it as an algorithm.
@@
0:  $\{(S_{n},X_{n},Y_{n})\}_{n=1}^{N},\{(S_{m}^{\star},X_{m}^{\star})\}_{m=1}^{M}$, Lipschitz constant $L$, confidence level $1-\alpha$, $\sigma^{2}$ (optional)
0:  A $(1-\alpha)$-confidence interval $I^{\Psi}$ for $\theta^{\star}_{\text{OLS},p}$
1:  $\Psi\leftarrow\text{1-NN}(\{S_{n}\}_{n=1}^{N},\{S_{m}^{\star}\}_{m=1}^{M})$ (Definition 5)
2:  $\hat{\theta}_{p}^{\Psi}\;\leftarrow\;e_{p}^{\top}\bigl{(}X^{\star\top}X^{\star}\bigr{)}^{-1}\,X^{\star\top}\,\Psi\,Y$
3:  $w\leftarrow e_{p}^{\top}\bigl{(}X^{\star\top}X^{\star}\bigr{)}^{-1}X^{\star\top},v^{\Psi}\leftarrow w\,\Psi$
4:  $B\leftarrow\!\sup_{g\in\mathcal{F}_{L}}\!\bigl{|}\,\sum_{m=1}^{M}w_{m}\,g(S^{\star}_{m})\!-\!\sum_{n=1}^{N}v^{\Psi}_{n}\,g(S_{n})\Bigr{|}.$ Compute with linear program (Section B.2).
5:  If $\sigma^{2}$ unknown, solve for the following estimator via quadratic program (Section B.5):
@@
8:  $I^{\Psi}\;\leftarrow\;\Bigl{[}\,\hat{\theta}_{p}^{\Psi}-B-c\,\Delta(\alpha)\;,\;\hat{\theta}_{p}^{\Psi}+B+c\,\Delta(\alpha)\Bigr{]}.$
@@
$\Psi_{mn}$ =$=\begin{cases}\frac{1}{K}&S_{n}\in\{K\text{ \small closest sourcelocations to }S^{\star}_{m}\}\\ 0&\text{ \small otherwise}\end{cases}.$ (11)
@@
When using a K-nearest neighbor matrix, there is an inherent bias–variance trade-off in choosing K. Increasing K broadens the geographic range of source observations used, hence the bias will increase; however, it also lowers the variance by averaging over more responses. We do not investigate this trade-off in detail. In the experiments we ran, we found $1$-nearest neighbor to work well. This is consistent with results in the mean estimation literature suggesting that if source and target distributions overlap substantially, the variance term remains manageable, even when using $1$-nearest neighbor approaches [47].
@@ B.2 Implementation of the Wasserstein Bound Calculation
We show in Section E.3 how Eq. 7 reduces to computing a Wasserstein-1 distance between empirical measures. To implement the Wasserstein distance calculation, we rely on the Python Optimal Transport library [16]. For the simulated experiments, we compute the cost matrix using Euclidean distances between spatial locations. For the real-world experiment, we use the Haversine distance to account for Earth’s curvature.
B.3 Use of Confidence Interval Lemma 6
In all experiments, we construct confidence intervals as in Lemma 6. A simpler alternative, which guarantees $1-\alpha$ coverage for all Gaussian distributions $\mathcal{N}(b,c^{2})$ with $b\in[-B,B]$, is to form two-sided confidence intervals for each $b\in[-B,B]$ and then take their union. This produces an interval of the form
@@ B.3 Use of Confidence Interval Lemma 6
However, the confidence intervals from Lemma 6 are never longer than this union-based approach, and the root-finding step required to compute them adds negligible overhead. Consequently, we opt for the intervals in Lemma 6 in all our experiments.
B.4 Computation of $\Delta$ for Lemma 6
Define $g(\Delta)=\Phi(\Delta)-\Phi(-\frac{2B}{c}-\Delta)-1+\alpha$. Our goal is to find a root of $g$ in the interval $[\Phi^{-1}(1-\alpha),\Phi^{-1}(1-\alpha/2)]$. We first show that $g$ is monotonic, that a root exists in this interval, and that the root is unique.
@@ B.4 Computation of $\Delta$ for Lemma 6
B.5 Computation of $\hat{\sigma}^{2}_{N}$ via quadratic programming
@@
The first obstacle is that the infimum is taken over an infinite-dimensional space. However, by the Kirszbraun theorem [29, 63], every $L$-Lipschitz function defined on a subset of $\mathbb{R}^{D}$ or $\mathbb{S}^{D}$ can be extended to an $L$-Lipschitz function on the whole domain. Since our objective function depends only on the values of $g$ at the source spatial locations, we only need to enforce the Lipschitz condition between all pairs of source spatial locations, not over the entire domain. Enforcing the Lipschitz condition at these $N$ source locations amounts to $N(N-1)/2$ linear inequality constraints.
@@
where $L$ is the Lipschitz constant, $\Gamma\in\mathbb{R}^{N^{2}-N}$ is the matrix of pairwise distances between distinct points in $S$, and $A\in\mathbb{R}^{(N^{2}-N)\times N}$ is a sparse matrix with exactly one $1$ and one $-1$ in distinct rows of each column, representing all such pairs in its rows.
@@
B.6 Scalable Estimation of $\sigma^{2}$
The quadratic programming approach for estimating $\sigma^{2}$ outlined in the main text (Eq. 10) and described in detail in the previous section does not scale well to large numbers of source locations. Therefore, for our synthetic experiment with $N=10{,}000$, we take a different approach.
@@ B.6 Scalable Estimation of $\sigma^{2}$
B.7 Implementation of Baseline Methods
@@
Ordinary Least Squares. We use the ordinary least squares implementation from statsmodel [55]. We use the default implementation, which calculates the variance as the average sum of squared residuals with a degrees-of-freedom correction, as in Eq. 2. We use a $t$-statistic to compute the corresponding confidence interval, which is again the default in statsmodel [55].
@@
We first optimize the maximum likelihood of a Gaussian process regression model with Matérn 3/2 kernel defined on the spatial source locations summed with a linear kernel defined on the covariates. This has the same likelihood as having a linear mean function in the covariates with a Gaussian prior over the weights, see Rasmussen and Williams, [49, Page 28]. We use the GPFlow [37] implementation of the likelihood, and L-BFGS [33] for numerical optimization of the likelihood. The optimization is initialized using the GPFlow default parameters for the mean and covariance functions. Once we have calculated the maximum likelihood parameters, we compute the posterior credible interval for $\theta$. The posterior over $\theta$ is Gaussian and has the closed form,
@@
Appendix C Simulations Additional Details

### v3

@@ 5 Discussion
We propose a new method for constructing confidence intervals for spatial linear models. We show via theory and experiments that our intervals accurately quantify uncertainty under misspecification and covariate shift. In experiments, our method is the only method that consistently achieves (or even comes close to) nominal coverage. We observe that, very commonly in spatial data analyses, covariates and responses may be observed at different, nonrandom locations in space. Since our method does not actually use the source covariate values in inference for $\theta^{\star}_{\text{OLS}}$, it can be applied in this common missing-data scenario. Though it requires additional work, we believe the ideas here will extend naturally to the widely used class of generalized linear models.
@@ Acknowledgments
This work was supported in part by a Social and Ethical Responsibilities of Computing (SERC) seed grant, an Office of Naval Research Early Career Grant, Generali, a Microsoft Trustworthy AI Grant, and NSF grant 2214177.
@@ Appendix A Extended Related Work
Bias in Spatial Regression. Linear models with a Gaussian process random effect — that is, models of the form
@@
with $g$ a Gaussian process and $\epsilon$ independent and identically distributed noise — are a classic tool in spatial regression and remain widely used in applications [75, 23, 26]. Whether to treat the covariates as fixed functions of spatial location or random is a topic of significant debate. In order for the model to be identifiable, it is generally necessary for the covariates to be thought of as random [21, Proposition 1]. However, Paciorek, [45] observed it may be more reasonable to think of covariates as fixed (the perspective we also take). If $X$ and $g$ are not independent (in the random case) or close to orthogonal in the fixed case, Paciorek, [45] illustrates that bias is introduced into the estimation of $\theta$. And that the degree of this bias depends on the relative scales over which $X$ and $g$ vary. He gives a closed form for this bias under strong linear-Gaussian assumptions that depend on parameters that would generally be unknown. Page et al., [46] builds on the result of Paciorek, [45], and shows that bias in predictions made by the spatial model due to confounding may be small, despite biases in parameter estimation. Hodges and Reich, [29] proposed using restricted spatial regression, essentially explaining as much of the variation in the response as possible by the covariates to reduce bias in estimation due to spatially-correlated error. We take this approach in the GLS baseline used in our experiments. Nobre et al., [44] extended earlier work about bias due to spatial confounding by considering the case when multiple observations are available at the same spatial location with independent noise. They showed that spatial regression models can still be biased with repeated observations due to confounding.
In our work, we focus on bias in estimation of the target conditional OLS. Because we do not assume the model is well-specified, there is no ‘true’ value of the parameter to estimate. Instead, our goal is to estimate the ‘best’ (in a least squares sense) linear approximation to the response. We focus on the case when $X$ is a fixed function of location (Assumption˜1) and make weak non-parametric, smoothness assumptions (Assumption˜4). This contrasts with prior works that generally assume the covariates are random and make linear-Gaussian assumptions to calculate the bias [45, 46]. Our work shows how to directly incorporate bias into confidence intervals for $\theta$ under our weak assumptions, but does not address issues of (non-)identifiability in the estimation problem.
@@
Local Versus Global Regression Estimates. An alternative perspective on regression in spatial settings, referred to as geographically weighted regression (GWR), focuses on estimating the association between the response and covariates at each location in space [6, 19]. For this regression problem to be well-defined, the covariates at each spatial location must be treated as random variables. Like our method, this approach uses nearest neighbors or local smoothing approaches to estimate the local coefficients at each location in space. And — as in our work — because of this “data-borrowing” from nearby locations, bias is introduced into the estimation of these local regression coefficients. The confidence intervals reported for GWR typically do not account for this bias [79]. While Yu et al., [79] provide a formula for the bias introduced by “data-borrowing”, their formula relies on the assumption that the model is well-specified and depends on unobservable parameters.
We focus on the estimation bias in the estimation of global parameters. Although local parameter estimates are often of interest in spatial analyses [19, Chapter 1], global estimates are also useful as summary statistics for making decisions that impact regions or multiple localities. And we characterize the bias in these estimates under weak, non-parametric assumptions and incorporate it into our confidence intervals. Extensions of our approach to GWR coefficients are a promising direction for future research. However, this extension is not simple, as it involves accounting for randomness in the covariates.
@@
In contrast, our approach ensures nominal coverage even when de-biasing the estimate via importance weighting is not possible. This advantage comes at the cost of an additional regularity assumption (Assumption˜4), as well as often empirically wider confidence intervals.
Comparison to Semiparametric Inference for Partially Linear Model. Partially linear models take the form
$\mathbb{E}\left[Y|X,S|=\beta^{T}X+g(S)\mathbb{E}[X\mid S]=\chi(S)\right$.
where $\beta$ is the parameter of interest, $S$ is a nuisance (or control) variable, $X$ is the covariates, and $g$ is an unknown and possibly complicated function. These models are widely studied in the semiparametric literature. Among many others, Robins et al., [54], Robinson, [55], Chernozhukov et al., [11] focus on estimation of $\beta$, under the assumption that the triples ($S_{n},X_{n},Y_{n}$) are independent and identically distributed across data indices $n$.
In many spatial applications, it isn’t reasonable to think of the nuisance variable (geographic space) as sampled independently and identically, or even at regularly spaced locations. Observational data are often collected in a highly non-uniform way — densely in some regions, sparsely or not at all in others — due to physical constraints, accessibility, or policy decisions. This non-uniformity introduces distribution shifts when attempting to generalize inferences from one region to another. In the present work, we focus on inference with fixed spatial locations and do not impose regularity conditions on the sampling design. This setup allows us to quantify uncertainty in associations in cases where extrapolation to poorly-sampled geographic areas is required, or in cases with heavily clustered training locations and more uniform target locations.
A notable exception to the assumption of fully i.i.d. data in the semiparametric literature is Heckman, [27]. Heckman, [27] considered time as a nuisance variable, and assumed this nuisance variable was one-dimensional and sampled densely and in a sufficiently regular way. In contrast, we allow for multiple spatial dimensions and do not require regularity assumptions about the sampling design.
Comparison to KNN Imputation. Our estimator is closely related to KNN imputation [66]. The $1$-Nearest Neighbor point estimate we consider is equivalent to 11NN imputation for the response, but using only the spatial coordinates — not the covariate values — as features when performing imputation. Directly applying KNN imputation to fill in the target response and then calculating standard confidence intervals would not lead to correct estimates of the variance because of repeated use of training data when imputing missing values — we calculate the variances accounting for potential repeated use of these responses. Additionally, using our smoothness assumptions, we are able to quantify additional uncertainty due to potential bias introduced when imputing the missing response values at the target locations. Because our confidence intervals account for potential bias and calculate the variance accounting for the weight assigned to each training example, our confidence intervals are guaranteed to be conservative, whereas confidence intervals calculated after data imputation need not be.
@@ Appendix B Implementation Details
In this section, we describe the implementation details of our method. Particularly, this involves computing the upper bound on the bias, computation of $\Delta$ in Lemma˜6 and computation of $\hat{\sigma}^{2}_{N}$ from Eq.˜10. To summarize our method, we state it as an algorithm.
@@
$\{(S_{n},X_{n},Y_{n})\}_{n=1}^{N},\{(S_{m}^{\star},X_{m}^{\star})\}_{m=1}^{M}$, Lipschitz constant $L$, confidence level $1-\alpha$, $\sigma^{2}$(optional) A $(1-\alpha)$-confidence interval $I^{\Psi}$for $\theta^{\star}_{\text{OLS},p}$
[…]
5: If $\sigma^{2}$ unknown, solve for the following estimator via quadratic program (Section˜B.7):
@@
7:Find $\Delta(\alpha)$ satisfying $\Phi\bigl(\Delta(\alpha)\bigr)\;-\;\Phi\Bigl(-2B/c-\Delta(\alpha)\Bigr)\;=\;1-\alpha,$ by root finding algorithm (Section˜B.6).
[…]
@@
$\Psi_{mn}$ $=\begin{cases}\frac{1}{K}&S_{n}\in\{K\text{ \small closest source locations to }S^{\star}_{m}\}\\ 0&\text{ \small otherwise}\end{cases}.$ (11)
@@
When using a K-nearest neighbor matrix, there is an inherent bias–variance trade-off in choosing K. Increasing K broadens the geographic range of source observations used, hence the bias will increase; however, it also lowers the variance by averaging over more responses. The degree to which increasing $K$ introduces additional bias depends on the density of data relative to the Lipschitz constant — the benefit of smoother weighting schemes (e.g. $K>1$) generally becomes more pronounced as the density of spatial sampling increases since smoother weighting schemes naturally leverage multiple nearby points, reducing variance while controlling bias effectively. We do not investigate this trade-off in detail. In the experiments we ran, we found $1$-nearest neighbor to work well. This is consistent with results in the mean estimation literature suggesting that if source and target distributions overlap substantially, the variance term remains manageable, even when using $1$-nearest neighbor approaches [50].
@@ B.2 Implementation of the Wasserstein Bound Calculation
We show in Section˜F.3 how Eq.˜7 reduces to computing a Wasserstein-1 distance between empirical measures. To implement the Wasserstein distance calculation, we rely on the Python Optimal Transport library [18]. For the simulated experiments, we compute the cost matrix using Euclidean distances between spatial locations. For the real-world experiment, we use the Haversine distance to account for Earth’s curvature.
B.3 Choice of Lipschitz Constant Based on Prior Research: Example for Air Pollution
[…]
B.4 Confidence-Interval Widths
In some experiments, most notably the tree-cover analysis in Fig.˜3, our confidence intervals can appear wide. This increased width arises naturally from our explicit control of bias under extrapolation: when test locations differ substantially from training locations, the Wasserstein-based bias bound enlarges, yielding intervals that faithfully represent genuine uncertainty rather than methodological conservatism. By contrast, existing methods (e.g., OLS, sandwich, and importance-weighted approaches) produce much narrower intervals but fail to achieve nominal coverage, often approaching zero coverage in Fig.˜1 and Fig.˜2. Our method therefore yields the narrowest intervals among those that do maintain validity. Producing still-narrower intervals would be trivial if one were willing to sacrifice coverage—after all: a zero-width interval achieves minimal size but no inferential meaning. Finally, we note that our intervals are typically informative in practice: in many settings they remain sufficiently tight to exclude zero, providing clear conclusions about the sign and magnitude of associations.
In practice, interval width also reflects the spatial density of available training data. When observations are closely spaced relative to the smoothness scale implied by the Lipschitz constant, the bias bound remains small and the resulting intervals are narrow. As the spacing grows, the bound naturally increases, producing wider intervals that correctly reflect the added uncertainty from extrapolation. Thus, even when data are sparse, the width of our intervals provides a meaningful indication of the reliability of inferences under the assumed smoothness level.
B.5 Use of Confidence Interval Lemma 6
In all experiments, we construct confidence intervals as in Lemma˜6. A simpler alternative, which guarantees $1-\alpha$ coverage for all Gaussian distributions $\mathcal{N}(b,c^{2})$ with $b\in[-B,B]$, is to form two-sided confidence intervals for each $b\in[-B,B]$ and then take their union. This produces an interval of the form
@@ B.5 Use of Confidence Interval Lemma 6
However, the confidence intervals from Lemma˜6 are never longer than this union-based approach, and the root-finding step required to compute them adds negligible overhead. Consequently, we opt for the intervals in Lemma˜6 in all our experiments.
B.6 Computation of $\Delta$ for Lemma 6
Define $g(\Delta)=\Phi(\Delta)-\Phi(-\frac{2B}{c}-\Delta)-1+\alpha$. Our goal is to find a root of $g$ in the interval $\Phi^{-1}(1-\alpha),\Phi^{-1}(1-\alpha/2)]$. We first show that $g$ is monotonic, that a root exists in this interval, and that the root is unique.
@@ B.6 Computation of $\Delta$ for Lemma 6
B.7 Computation of $\hat{\sigma}^{2}_{N}$ via quadratic programming
@@
The first obstacle is that the infimum is taken over an infinite-dimensional space. However, by the Kirszbraun theorem [32, 69], every $L$-Lipschitz function defined on a subset of $\mathbb{R}^{D}$ or $\mathbb{S}^{D}$ can be extended to an $L$-Lipschitz function on the whole domain. Since our objective function depends only on the values of $g$ at the source spatial locations, we only need to enforce the Lipschitz condition between all pairs of source spatial locations, not over the entire domain. Enforcing the Lipschitz condition at these $N$ source locations amounts to N(N1)/2N(N-1)/2 linear inequality constraints.
@@
where $L$ is the Lipschitz constant, $\Gamma\in\mathbb{R}^{N^{2}-N}$ is the matrix of pairwise distances between distinct points in $S$, and $A\in\mathbb{R}^{(N^{2}-N)\times N}$ is a sparse matrix with exactly one $1$ and one 1-1 in distinct rows of each column, representing all such pairs in its rows.
@@
B.8 Scalable Estimation of $\sigma^{2}$
The quadratic programming approach for estimating $\sigma^{2}$ outlined in the main text (Eq.˜10) and described in detail in the previous section does not scale well to large numbers of source locations. Therefore, for our synthetic experiment with N=10, [^0] N=10{,}000, we take a different approach.
@@ B.8 Scalable Estimation of $\sigma^{2}$
B.9 Implementation of Baseline Methods
@@
Ordinary Least Squares. We use the ordinary least squares implementation from statsmodel [60]. We use the default implementation, which calculates the variance as the average sum of squared residuals with a degrees-of-freedom correction, as in Eq.˜2. We use a $t$-statistic to compute the corresponding confidence interval, which is again the default in statsmodel [60].
@@
We first optimize the maximum likelihood of a Gaussian process regression model with Matérn 3/2 kernel defined on the spatial source locations summed with a linear kernel defined on the covariates. This has the same likelihood as having a linear mean function in the covariates with a Gaussian prior over the weights, see Rasmussen and Williams, [52, Page 28]. We use the GPFlow [40] implementation of the likelihood, and L-BFGS [36] for numerical optimization of the likelihood. The optimization is initialized using the GPFlow default parameters for the mean and covariance functions. Once we have calculated the maximum likelihood parameters, we compute the posterior credible interval for $\theta$. The posterior over $\theta$ is Gaussian and has the closed form,
@@
Appendix C Extension To Multivariate Responses
We focus on a scalar response $Y_{n}\in\mathbb{R}$. We expect similar machinery can be adapted to confidence intervals for each coordinate of $Y_{n}\in\mathbb{R}^{D}$. Namely, we can treat each component $Y^{(d)}_{n}\in\mathbb{R}^{D}$ as a separate univariate problem, and apply our Lipschitz‐based bias bound and variance calibration to each coordinate. If one desires simultaneous coverage over all $D$ outputs, a straightforward Bonferroni correction (i.e. defining $\alpha^{\prime}=\alpha/D$ and applying our method to construct $1-\alpha^{\prime}$ confidence intervals for each coordinate) or another family‐wise error control can be used. An exploration of improvements of this Bonferroni correction approach for multivariate responses would be an interesting direction for future work.
Appendix D Simulations Additional Details